OpenStudy (anonymous):

y+xcosy=(x^2)y f'X=?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

implicit derivative i assume? or is this partial derivative?

OpenStudy (anonymous):

I have no idea :/

OpenStudy (amistre64):

i have to wonder about the notation used in the post. \(f_x\) generally denotes a partial derivative

OpenStudy (anonymous):

its f(x)

OpenStudy (amistre64):

if you simply mean y' then its implicit

OpenStudy (amistre64):

for starters, forget everything you think you know about how x and y act .... thats useless ideas. y+xcosy=(x^2)y y is a power rule; how would you take the derivative of y? \[\frac d{dx}(y)=?\]

OpenStudy (anonymous):

f(x)=x^4tan^(−1)(3x) this

OpenStudy (anonymous):

question sorry

OpenStudy (anonymous):

Implicit Differentiation

OpenStudy (amistre64):

... so f'(x) of that inverse tangent function?

OpenStudy (amistre64):

recall the product rule: \(D_x(fg)=f'g+fg'\) correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if we let f=x, what is f' equal?

OpenStudy (amistre64):

prolly used a bad choice of letters for the product rule, but that can be overlooked :)

OpenStudy (anonymous):

what is the first step to solve this

OpenStudy (amistre64):

the first step is in getting the correct equation; is it: \[f(x)=x^{4tan^{-1}(3x)}\] OR \[f(x)=x^{4}~{tan^{-1}(3x)}\]

OpenStudy (anonymous):

the second equation

OpenStudy (amistre64):

then that is simply a product of the terms: x^4 and: tan^(-1)(3x) you know the power rule, and the second term is gonna employ the chain rule as well

OpenStudy (anonymous):

4x^3? for the power and I am stuck with chain rules. I have veen practing it but not improving :/

OpenStudy (amistre64):

if we want to follow a systematic setup then maybe do this: \[D_x(uv)=u'v+uv'\] \[u=x^4~:~u'=? \] \[v=tan^{-1}3x~:~v'=?\]

OpenStudy (anonymous):

u=4x^3

OpenStudy (amistre64):

u' = 4x^3 good \[D_x(uv)=u'v+uv'\] \[D_x(uv)=4x^3~tan^{-1}(3x)+x^4v'\]

OpenStudy (anonymous):

v' is the hard one

OpenStudy (amistre64):

the chain rule is such that we simply peel away the functions to form a product\[D_x(f(g(h(x))))=f'(g(h(x)))~~g'(h(x))~~h'(x)~~x'\]

OpenStudy (amistre64):

\[D_x[tan^{-1}(3x)]=(tan^{-1})'(3x)~~(3x')\] \[D_x[tan^{-1}(3x)]=3(tan^{-1})'(3x)\] ill have to review for myself what the tan-1' is :)

OpenStudy (anonymous):

ohh haha this is why they call it chain rules! it looks like a chain

OpenStudy (amistre64):

let: y=arctan(x) tan(y)=x sec^2(y) y' = 1 y' = 1/(sec^2(y)) ; and we know y=arctan(x) y' = 1/(sec^2(arctan(x))  yeah, 1/(1+x^2)